OpenStudy (anonymous):

Solve. 6x=9+x^2 a.3 b.-3 c.-3,3 d. no solution

jimthompson5910 (jim_thompson5910):

6x=9+x^2 0=9+x^2-6x 0 = x^2 - 6x + 9 x^2 - 6x + 9 = 0 (x-3)^2 = 0 x - 3 = 0 x = 3 So the only solution is x = 3

OpenStudy (lgbasallote):

you can try to plug in each of your choices into x